OpenStudy (anonymous):

solve the equation |k-10|=3

OpenStudy (kirbykirby):

if (k-10) < 0, then -(k-10) = 3 -k+10 = 3 -k = -7 k = 7 if (k-10) > 0, then k-10 = 3 k = 13

OpenStudy (kirbykirby):

This is from the property of absolute value: |x| = -x if x < 0 = x if x > 0 here, just think that (k-10) is your x, so substitute it to get |k-10|...